aldavis:

The point P(6, –1) is translated according to the rule (x, y) → (x – 5, y – 2). What are the coordinates of P'?

sophia333:

(6-5=1, -1-2=-3)

sophia333:

So do you know your answer?

sophia333:

@aldavis are you there?

XioGonz:

(6,-1) to ---> (x-5,y-2) So that we get (6-5,-1-2) (when we plug in the values of x and y-axis) And we get (1,-1+-2) =(1,-3)